OpenStudy (tia_sweetheart17):

factor completely. 3x^2+22x+40

OpenStudy (anonymous):

3x^2 + 22x + 40 = 0 3x^2 + 12x + 10x + 40 = 0 3x (x + 4) +10 (x + 4) = 0 (3x + 10)(x + 4) = 0 3x + 10 = 0 => x = -10/3 x + 4 = 0 => x = -4